OpenStudy (anonymous):

complete the standard form of the eqaution -5y=10-2x. ?x +? y=10

Directrix (directrix):

-5y=10-2x The 10 has to be on the right side but the -2x has to move to the left side.

Directrix (directrix):

-5y = 10 - 2x 2x + 2x =================== 2x - 5y = 10

Directrix (directrix):

Read more here: Standard Form of a Linear Equation http://www.mathsisfun.com/algebra/standard-form.html